OpenStudy (anonymous):

Solve: −3x − 7 = 17 −3.3 4 8 −8

OpenStudy (anonymous):

alright, -3x-7=17 +7 to both sides to get x alone -3x=17+7...or....-3x=24 now divied both sides by -3 to get the answer x=-8